# Almost Surely Asymptotically Constant Graph Neural Networks

### v2

@@ D.1 Combinatorial tools for the sparse case
|$\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T,\bar{w},y)\}|}{n}\leq\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(v)\cong\operatorname{\mathcal{N}}^{T}_{\ell}(y)\}|}{n}$
By Lemma 5.6 we have that:
|$\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(v)\cong\operatorname{\mathcal{N}}^{T}_{\ell}(y)\}|}{n}$
converges to some limit $q_{T}$ asymptotically almost surely, while by Lemma D.6 we have that:
$\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u})|}{n}$
converges to 00, so the result follows. ∎
@@
We now begin the proof of Theorem D.1. Recall from the body that we will do this via Lemma 5.7, which requires us to define some Lipschitz controllers on neighbourhood types that approximate a given term $\pi$ relative to a neighbourhood type.
@@
Recall from Lemma 5.7 that our goal is to ensure the following correctness property for our controllers $e_{\pi}^{T}$.
@@
    Note that $\ell=0$. Let $\ell^{\prime}=\max(\operatorname{\mathrm{Reach}}(\rho),\operatorname{\mathrm{Reach}}(\eta))$. For any tuple $\bar{u}$ and $(k+1)$-rooted graph $(T^{\prime},\bar{u},v)$ let:
    $\llbracket T^{\prime}(\bar{u})\rrbracket\coloneq|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{w},y)\cong T^{\prime}\}|$
    By the parameterized Weak Local Convergence result, Theorem D.3, for every such $T^{\prime}$ there is $q_{T^{\prime}}\in[0,1]$ such that for all $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$, for every $k$-tuple $\bar{u}$ we have that:
    $\left|\frac{\llbracket T^{\prime}(\bar{u})\rrbracket}{n}-q_{T^{\prime}}\right|<\epsilon$
    Let $\mathcal{T}_{k+1}$ be the set of isomorphism types of $(k+1)$-rooted graphs. Note that:
    $\sum_{T^{\prime}\in\mathcal{T}_{k+1}}q_{T^{\prime}}\leq 1$
    Indeed, if this sum were greater than $1$, there would be some finite set $S\subseteq\mathcal{T}_{k+1}$ such that $\sum_{T^{\prime}\in S}q_{T^{\prime}}>1$, and so by making $\epsilon$ small enough we would have that the proportion of nodes in the graph which satisfy some $T^{\prime}\in S$ is greater than $1$, which is impossible.
    Consider any $(k+1)$-rooted graph $(T^{\prime},\bar{w},y)$. In order to define the controller, we need to identify the nodes in the canonical enumeration corresponding to $\operatorname{\mathcal{N}}_{\ell^{\prime}-1}^{T^{\prime}}(\bar{u})$. Let:
    $r_{T^{\prime}}\coloneqq\left|\operatorname{\mathcal{N}}_{\ell^{\prime}-1}^{T^{\prime}}(\bar{u})\right|$
    Given tuples of nodes $\bar{\bm{a}}=({\bm{a}}_{1},\ldots,{\bm{a}}_{r_{T^{\prime}}})$ and $\bar{\bm{b}}=({\bm{b}}_{r_{T^{\prime}}+1},\ldots,{\bm{b}}_{|T^{\prime}|})$, let $\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}})$ be the $|T^{\prime}|$-tuple of vectors obtained by assigning the ${\bm{a}}_{i}$’s to the nodes in $\operatorname{\mathcal{N}}_{\ell^{\prime}-1}^{T^{\prime}}(\bar{u})$ and the ${\bm{b}}_{i}$’s to the remaining nodes in $T^{\prime}$, according to the canonical enumeration of $T^{\prime}$.
@@
    $f_{\pi}^{T}(\bar{\bm{a}})\coloneqq\sum_{T^{\prime}\in\mathcal{T}_{k+1}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[e_{\rho}^{T^{\prime}}(\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}}))\cdot h(e_{\eta}^{T^{\prime}}(\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}})))\right]\cdot q_{T^{\prime}}$
@@
    $g_{\pi}^{T}({\bm{a}}_{1},\ldots,{\bm{a}}_{k})\coloneqq\sum_{T^{\prime}\in\mathcal{T}_{k+1}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[h(e_{\eta}^{T^{\prime}}(\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}})))\right]\cdot q_{T^{\prime}}$
@@
    We first argue that the sum in the definition is well-defined. Since:
@@
    are bounded, and $\sum_{T^{\prime}\in\mathcal{T}_{k+1}}q_{T^{\prime}}$ converges, these sums converges absolutely.
    Now, take $\epsilon,\theta>0$. Take a finite $S\subseteq\mathcal{T}_{k+1}$ such that:
@@
    and each $q_{T^{\prime}}>0$ for $T^{\prime}\in S$.
    The guarantee given when we applied the parameterized version of Weak Local Convergence, Theorem D.3, is that there is $N_{1}$ such that for all $n\geq N_{1}$ with probability at least $1-\theta$, for all $T^{\prime}\in S$ and every $k$-tuple $\bar{u}$ we have that:
    $\left|\frac{\llbracket T^{\prime}(\bar{u})\rrbracket}{n}-q_{T^{\prime}}\right|<\epsilon$ ($\ddagger$)
@@
    Note that $f_{\pi}^{\circ}$ is bounded, say by $\Lambda$. Then for every $\gamma>0$, by Hoeffding’s Inequality (Theorem B.4) for any tuple $\bar{u}$ and $T^{\prime}\in S$ we have that:
@@
    $\left\lVert\frac{1}{n}\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}f_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})-q_{T^{\prime}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[f_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]\right\rVert<\epsilon$ ($\triangle_{f}$)
    Similarly, there is $N_{5}$ such that for all $n\geq N_{5}$ with probability at least $1-\theta$, for all $\bar{u}$ and $T^{\prime}$ we have:
    $\left\lVert\frac{1}{n}\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})-q_{T^{\prime}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]\right\rVert<\epsilon$ ($\triangle_{g}$)
@@
    Take any $k$-tuple of nodes $\bar{u}$. It suffices to show, using the definition of the interpretation of the weighted mean operator, that:
@@
    $\frac{1}{n}\sum_{v}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})=\frac{1}{n}\sum_{T^{\prime}\in\mathcal{T}_{k+1}}\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})$
@@
    $\left\lVert\frac{1}{n}\sum_{T^{\prime}\in S}\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})-\frac{1}{n}\sum_{T^{\prime}\in S}q_{T^{\prime}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]\right\rVert\leq\epsilon$
@@
    $\left\lVert\frac{1}{n}\sum_{T^{\prime}\in S}q_{T^{\prime}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]-g_{\pi}^{T}(\bar{\bm{a}}^{\bar{u}})\right\rVert<\epsilon\Lambda$
This completes the proof of the inductive construction of the controllers, thus proving Lemma 5.7. ∎
@@ Appendix F Additional experiments
Figure 5: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT ane GPS+RW model initialization over the $\mathrm{ER}(n,p(n)=0.1)$ graph distributions, as we draw increasingly larger graphs.
Figure 6: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT ane GPS+RW model initialization over the $\mathrm{ER}(n,p(n)=\frac{\log{n}}{n})$ graph distributions, as we draw increasingly larger graphs.
Figure 7: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT ane GPS+RW model initialization over the $\mathrm{ER}(n,p(n)=\frac{1}{50n})$ graph distributions, as we draw increasingly larger graphs.
Figure 8: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT ane GPS+RW model initialization over the SBM graph distributions, as we draw increasingly larger graphs.
Figure 9: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT ane GPS+RW model initialization over the $\mathrm{BA}(n,m=5)$ graph distributions, as we draw increasingly larger graphs.

### v3

@@ D.1 Combinatorial tools for the sparse case
|$\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u},v)\cong(T^{\prime},\bar{w},y)\}|}{n}\leq\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(v)\cong\operatorname{\mathcal{N}}^{T^{\prime}}_{\ell}(y)\}|}{n}$
By Lemma 5.5 (with $k=1$) we have that:
|$\frac{|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell}(v)\cong\operatorname{\mathcal{N}}^{T^{\prime}}_{\ell}(y)\}|}{n}$
converges to some limit $q_{T\mid T^{\prime}}\coloneqq q_{\operatorname{\mathcal{N}}^{T^{\prime}}_{\ell}(y)}$ asymptotically almost surely, while:
$\frac{|\operatorname{\mathcal{N}}^{G}_{\ell}(\bar{u})|}{n}=\frac{|T|}{n}$
converges to 00. This completes the proof for Case 1.
Finally, to show that $\sum_{T^{\prime}\in\mathsf{Ext}(T)}q_{T^{\prime}\mid T}=1$, note the set:
$\{\operatorname{\mathcal{N}}^{T^{\prime}}_{\ell}(y)\mid(T^{\prime},\bar{w},y)\in\mathsf{Ext}(T^{\prime})\}$
contains all $1$-rooted graphs up to isomorphism which can be the $\ell$-neighbourhood of a node. Therefore, by the “moreover” part of Definition D.5 we have that:
$\sum_{T^{\prime}\in\mathsf{Ext}(T)}q_{T^{\prime}\mid T}=\sum_{T^{\prime}\in\mathsf{Ext}(T)}q_{\operatorname{\mathcal{N}}^{T^{\prime}}_{\ell}(y)}=1$
@@
We now begin the proof of Theorem D.1. Recall from the body that we will do this via Theorem 5.6, which requires us to define some Lipschitz controllers on neighbourhood types that approximate a given term $\pi$ relative to a neighbourhood type.
@@
Recall from Theorem 5.6 that our goal is to ensure the following correctness property for our controllers $e_{\pi}^{T}$.
@@
    Note that $\ell=0$. Let $\ell^{\prime}=\max(\operatorname{\mathrm{Reach}}(\rho),\operatorname{\mathrm{Reach}}(\eta))$. Take a $k$-rooted graph $(T,\bar{w})$. For any tuple $\bar{u}$ and $(k+1)$-rooted graph $(T^{\prime},\bar{w},y)$ extending $(T,\bar{w})$ let:
    $\llbracket T^{\prime}(\bar{u})\rrbracket\coloneq|\{v\mid\operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong(T^{\prime},\bar{w},y)\}|$
    By the parameterized Weak Local Convergence result, Theorem D.4, for every such $T^{\prime}$ extending there is a relative neighborhood weight $q_{T\mid T^{\prime}}\in[0,1]$ such that for all $\epsilon,\theta>0$ there is $N\in\mathbb{N}$ such that for all $n\geq N$ with probability at least $1-\theta$, for every $k$-tuple $\bar{u}$ such that $\operatorname{\mathcal{N}}_{\ell}(\bar{u})=T$ we have that:
    $\left|\frac{\llbracket T^{\prime}(\bar{u})\rrbracket}{n}-q_{T\mid T^{\prime}}\right|<\epsilon$
    and moreover the relative neighborhood weights sum to one:
    $\sum_{T^{\prime}\in\mathsf{Ext}(T)}q_{T\mid T^{\prime}}=1$
    Consider any $(T^{\prime},\bar{w},y)\in\mathsf{Ext}(T)$. In order to define the controller, we need to identify the nodes in the canonical enumeration corresponding to $\operatorname{\mathcal{N}}_{\ell^{\prime}-1}^{T}(\bar{u})$. Let:
    $r_{T}\coloneqq\left|\operatorname{\mathcal{N}}_{\ell^{\prime}-1}^{T}(\bar{u})\right|$
    Given tuples of $\mathbb{R}^{d}$-vectors $\bar{\bm{a}}=({\bm{a}}_{1},\ldots,{\bm{a}}_{r_{T}})$ and $\bar{\bm{b}}=({\bm{b}}_{r_{T}+1},\ldots,{\bm{b}}_{|T^{\prime}|})$, let $\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}})$ be the $|T^{\prime}|$-tuple of vectors obtained by assigning the ${\bm{a}}_{i}$’s to the nodes in $\operatorname{\mathcal{N}}_{\ell^{\prime}-1}^{T}(\bar{u})$ and the ${\bm{b}}_{i}$’s to the remaining nodes in $T^{\prime}$, according to the canonical enumeration of $T^{\prime}$.
@@
    $f_{\pi}^{T}(\bar{\bm{a}})\coloneqq\sum_{T^{\prime}\in\mathsf{Ext}(T)}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[e_{\rho}^{T^{\prime}}(\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}}))\cdot h(e_{\eta}^{T^{\prime}}(\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}})))\right]\cdot q_{T\mid T^{\prime}}$
@@
    $g_{\pi}^{T}({\bm{a}}_{1},\ldots,{\bm{a}}_{k})\coloneqq\sum_{T^{\prime}\in\mathsf{Ext}(T)}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[h(e_{\eta}^{T^{\prime}}(\operatorname{\mathrm{Arrange}}_{T^{\prime}}(\bar{\bm{a}},\bar{\bm{b}})))\right]\cdot q_{T\mid T^{\prime}}$
@@
    Since:
@@
    are bounded, and $\sum_{T^{\prime}\in\mathsf{Ext}(T)}q_{T\mid T^{\prime}}$ converges, these sums converges absolutely, and hence the controller is well-defined.
    Take $\epsilon,\theta>0$. Take a finite $S\subseteq\mathsf{Ext}(T)$ such that:
@@
    and each $q_{T\mid T^{\prime}}>0$ for $T^{\prime}\in S$.
    The guarantee given when we applied the parameterized version of Weak Local Convergence, Theorem D.4, is that there is $N_{1}$ such that for all $n\geq N_{1}$ with probability at least $1-\theta$, for all $T^{\prime}\in S$ and every $k$-tuple $\bar{u}$ such that $\operatorname{\mathcal{N}}_{\ell}(\bar{u})=T$ we have that:
    $\left|\frac{\llbracket T^{\prime}(\bar{u})\rrbracket}{n}-q_{T\mid T^{\prime}}\right|<\epsilon$ ($\ddagger$)
@@
    Note that $f_{\pi}^{\circ}$ is bounded, say by $\Lambda$. Then for every $\gamma>0$, by Hoeffding’s Inequality (Theorem B.4) for any tuple $\bar{u}$ such that $\operatorname{\mathcal{N}}_{\ell}(\bar{u})=T$ and $T^{\prime}\in S$ we have that:
@@
    $\left\lVert\frac{1}{n}\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}f_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})-q_{T\mid T^{\prime}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[f_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]\right\rVert<\epsilon$ ($\triangle_{f}$)
    Similarly, there is $N_{5}$ such that for all $n\geq N_{5}$ with probability at least $1-\theta$, for all $\bar{u}$ such that $\operatorname{\mathcal{N}}_{\ell}(\bar{u})=T$ and $T^{\prime}\in\mathsf{Ext}(T)$ we have:
    $\left\lVert\frac{1}{n}\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})-q_{T\mid T^{\prime}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]\right\rVert<\epsilon$ ($\triangle_{g}$)
@@
    Take any $k$-tuple of nodes $\bar{u}$ such that $\operatorname{\mathcal{N}}_{\ell}(\bar{u})=T$. It suffices to show, using the definition of the interpretation of the weighted mean operator, that:
@@
    $\frac{1}{n}\sum_{v}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})=\frac{1}{n}\sum_{T^{\prime}\in\mathsf{Ext}(T)}\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})$
@@
    $\left\lVert\frac{1}{n}\sum_{T^{\prime}\in S}\sum_{\begin{subarray}{c}v\\ \operatorname{\mathcal{N}}^{G}_{\ell^{\prime}}(\bar{u},v)\cong T^{\prime}\end{subarray}}g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}}^{(\bar{u},v)})-\frac{1}{n}\sum_{T^{\prime}\in S}q_{T\mid T^{\prime}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]\right\rVert\leq\epsilon$
@@
    $\left\lVert\frac{1}{n}\sum_{T^{\prime}\in S}q_{T\mid T^{\prime}}\operatorname*{\mathbb{E}}_{\bar{\bm{b}}\sim\mathbb{F}}\left[g_{\pi}^{\circ}(\bar{\bm{a}}^{\bar{u}},\bar{\bm{b}})\right]-g_{\pi}^{T}(\bar{\bm{a}}^{\bar{u}})\right\rVert<\epsilon\Lambda$
This completes the proof of the inductive construction of the controllers, thus proving Theorem 5.6. ∎
@@ Appendix F Additional experiments
Figure 6: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT and GPS+RW model initialization over the $\mathrm{ER}(n,p(n)=0.1)$ graph distributions, as we draw increasingly larger graphs.
Figure 7: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT and GPS+RW model initialization over the $\mathrm{ER}(n,p(n)=\frac{\log{n}}{n})$ graph distributions, as we draw increasingly larger graphs.
Figure 8: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT and GPS+RW model initialization over the $\mathrm{ER}(n,p(n)=\frac{1}{50n})$ graph distributions, as we draw increasingly larger graphs.
Figure 9: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT and GPS+RW model initialization over the SBM graph distributions, as we draw increasingly larger graphs.
Figure 10: The 5 class probabilities (in different colours) of a MeanGNN, GCN, GAT and GPS+RW model initialization over the $\mathrm{BA}(n,m=5)$ graph distributions, as we draw increasingly larger graphs.
Figure 11: A three-layer GCN with hidden dimension 128 is trained on the ENZYMES dataset with one class removed so that there are five output classes. This model is then run on graphs drawn from $\mathrm{ER}(n,p(n)=0.1)$ for increasing sizes $n$, and the mean output probabilities are recorded, along with standard deviation
Appendix G Acknowledgements
This research was funded in whole or in part by EPSRC grant EP/T022124/1. For the purpose of Open Access, the authors have applied a CC BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.